OpenStudy (anonymous):

solve for x 20x^2-7x=6

OpenStudy (anonymous):

\[20x^2-7x=6 \rightarrow 20x^2-7x-6=0\rightarrow(4x-3)(5x+2)=0\rightarrow x=\frac{ 3 }{ 4 } ; x=\frac{ -2 }{ 5 }\]